The ratio of the measures of three sides of a triangle are 1/2: 1/3: 1/5, and its perimeter is 6.2 centimeters. Find the measure of each side of the triangle.
Answer by rfer(16322)

15/30x+10/30x+6/30x=6.2
31/30x=6.2
x=6.2*30/31
x=186/31
x=6
------------
1/2x=3
1/3x=2
1/5x=1.2